OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x + 2y = 16 y = x - 4

OpenStudy (anonymous):

The second equation you have is \(y=x-4\) so if you multiply by 2 you get that \(2y=2x-8\) now fill in/substitute this into the first equation to get \[2x+2y=2x +(2x-8)=16\] in other words \(4x-8 = 16\) or rearranging we get \(4x=24\) so \(x=6\). Now you know \(x\) you can fill it back into either of the given equations to find \(y\)